OpenStudy (anonymous):

Write the equation of a line in slope-intercept form that has a slope of -1/4 and passes through the point (8, -1). Slope should be a reduced improper fraction.

OpenStudy (mathstudent55):

The slope-intercept form of the equation of a line is: \(y = mx + b\) where \(m\) = slope, and \(b\) = the y-intercept

OpenStudy (mathstudent55):

You are told that the slope is \( -\dfrac{1}{4} \), so you can already substitute that value into m in the equation: \( y = -\dfrac{1}{4}x + b \)

OpenStudy (anonymous):

okay so I would set it up something like this 

OpenStudy (mathstudent55):

Now you need to find \(b\). To do that, use the point you are given, (8, -1), and replace \(x\) with 8 and \(y\) with -1 in the equation. Then solve for \(b\): \(y = -\dfrac{1}{4}x + b\) \(-1 = -\dfrac{1}{4}(8) + b\) \(-1 = -2 + b\) \(1 = b\) Now that we know \(b = 1\), just substitute 1 for b in our equation: \(y = -\dfrac{1}{4}x + 1\)

OpenStudy (anonymous):

but the 1 is a negative isnt it?

OpenStudy (mathstudent55):

No, it's positive.

OpenStudy (anonymous):

oh.

OpenStudy (mathstudent55):

Look at our result: \(y = -\dfrac{1}{4}x + 1 \) Just from looking at it, you can see the slope is \( -\dfrac{1}{4} \). Now let's test point (8, -1): \(y = -\dfrac{1}{4}x + 1 \) \(-1 = -\dfrac{1}{4}(8) + 1 \) \(-1 = -2 + 1 \) \(-1 = -1 \) Since -1 = -1 is a true statement, that means (8, -1) is a point on the line. The line has the slope we need, and the point we were asked about is on the line, so the equation of the line is correct with b = 1.

OpenStudy (anonymous):

so that's the answer. b=1...?

OpenStudy (mathstudent55):

The answer is the equation: \(y = -\dfrac{1}{4}x + 1 \)

OpenStudy (anonymous):

But how would I put that into an improper fraction?

OpenStudy (mathstudent55):

The slope here happens to be \(-\dfrac{1}{4} \). It is reduced, but since it's not an improper fraction, it can't be written as one. Did you by any chance miscopy the problem? Did your problem by any chance ask to find the line that is perpendicular to a line with slope \(-\dfrac{1}{4} \)?

OpenStudy (anonymous):

uhm.. No, it didn't. So i guess it cant be reduced or made into an improper fraction, that would be the answer?

OpenStudy (anonymous):

oh, never mind I got the answer :3 thanks for the help

OpenStudy (mathstudent55):

It is reduced. It simply is not an improper fraction.

OpenStudy (mathstudent55):

You're welcome.